OpenStudy (anonymous):

3x-y=10 -x+y=-4 3,-7 -1,3 3,-1 3,1

OpenStudy (anonymous):

3x-y=10 y=3x-10 putting the value u would get : -x+3x-10=-4 2x-10=-4 2x=6 x=3 and so 9-y=10 y=-1 so the answer is 3,-1